If I have two and a half dozen coins that total 20 cents.
What are the coins?

Answer by jim_thompson5910(35256)

two and a half dozen coins = (2 dozen) + (half dozen) = 2*12 + (1/2)*12 = 24+6 = 30

two and a half dozen coins = 30 coins


It's not possible to have 20 cents if you have 30 coins.